OpenStudy (anonymous):

Sketch the curve y=ln(x) and find the tangent line to this curve at the point where the curve crosses the x-axis. How do i do this on the computer?

OpenStudy (campbell_st):

ok... so here is a site that will graph the curve so you can see where it cuts the x-axis https://www.desmos.com/calculator next you will need to find the 1st derivative... do you know how to do that..?

OpenStudy (anonymous):

no!?

OpenStudy (campbell_st):

ok... well I don't know how you will find the equation of the tangent.... at that point... if you don't know how to differentiate

OpenStudy (campbell_st):

at least you're able to identify the point... that has to be part marks

OpenStudy (arbershabani97):

you don't need to know derivatives when you learn logs etc., so if you want to know how to solve just find y=e^x and y=ln(x) is the inverse of it, and as @campbell_st said you can do it in desmos.com

OpenStudy (arbershabani97):

I have no idea what it is, even though I have learned about logs etc :P

OpenStudy (campbell_st):

well to find a tangent you need a point, which you have and a slope... which you don't have. By find the 1st derivative... you get the equation of the slope of the tangent. then you substitute the x value for the point to find the slope of the tangent at the given point... then you can go and use point slope formula to find the equation of the tangent...

OpenStudy (anonymous):

the background info that i was also given was that air pressure decays exponentially at 0.4% per each rise of 30m above sea level. Could i use this information?

OpenStudy (campbell_st):

thanks @arbershabani97 2 nice curves... how do you find the slope... of the tangent to the log curve...

OpenStudy (campbell_st):

well I have no clue on the question other what you have posted... the background information doesn't make a lot of sense...

OpenStudy (arbershabani97):

you don't have to y=e^x when x=-1 --- y=0.368 when x=0 ---- y=1 when x=1---- y=2.718 when x=2 ---- y=7.389 and you just reflect those points to find y=ln(x) replace x with y when x=0.368 ---- y=-1 when x=1 ----- y=0 when x=2.718 ---- y=1

OpenStudy (arbershabani97):

you don't have to find slope*

OpenStudy (campbell_st):

what you have is a refection of the curve \[y = e^x\] in the line y = x this is what it looks like

OpenStudy (campbell_st):

but it doesn't help you find the slope of the tangent...

OpenStudy (arbershabani97):

sorry, I don't know how to find the slope

OpenStudy (arbershabani97):

i thought you were asking only for the graph

OpenStudy (campbell_st):

you need to slope of the tangent to y = ln(x) at (1, 0) to find the equation of the tangent.

OpenStudy (campbell_st):

perhaps you could just use http://www.wolframalpha.com/ it will give you a solution with little understanding

OpenStudy (anonymous):

could this be right?